# Logarithmic regret in the ergodic Avellaneda–Stoikov market making model

### v1

@@ A.17. Proof of Theorem 27
$\begin{split}\lambda^{+}\psi^{\kappa,+}(q)e^{-\kappa^{\star}\psi^{\kappa,+}(q)}+\lambda^{-}\psi^{\kappa,-}(q)&e^{-\kappa^{\star}\psi^{\kappa,-}(q)}-\phi q^{2}=\\ \gamma(\kappa;\kappa^{\star})&+\lambda^{+}e^{-\kappa^{\star}\psi^{\kappa,+}}\big{(}\hat{v}^{\psi^{\kappa}}(q-1;\kappa^{\star})-\hat{v}^{\psi^{\kappa}}(q;\kappa^{\star})\big{)}\bm{1}_{q>\underline{q}}\\ &+\lambda^{-}e^{-\kappa^{\star}\psi^{\kappa,-}}\big{(}\hat{v}^{\psi^{\kappa}}(q+1;\kappa^{\star})-\hat{v}^{\psi^{\kappa}}(q;\kappa^{\star})\big{)}\bm{1}_{q<\bar{q}},\end{split}$
where we ignore the indicator functions in the first line since $\psi^{\kappa,\pm}(q)e^{-\kappa^{\star}\psi^{\kappa,\pm}(q)}=0$ for $q=\bar{q},\underline{q}$, respectively. Also, by definition, we know that
$f(t,q,\delta^{\pm};\kappa^{\star})=\lambda^{+}\delta^{+}e^{-\kappa^{\star}\delta^{+}}+\lambda^{-}\delta^{-}e^{-\kappa^{\star}\delta^{-}}-\phi q^{2}.$
@@
(57) $\begin{split}&\mathcal{R}^{\Psi}(T)=\gamma(\kappa^{\star})T-\mathbb{E}_{q}\Big{[}\int_{0}^{T}f(t,Q_{t}^{\psi^{\kappa_{t}};\kappa^{\star}},\psi^{\kappa_{t}};\kappa^{\star})\,\mathrm{d}t\Big{]}\\ &=\mathbb{E}\Big{[}\int_{0}^{T}\big{(}\gamma(\kappa^{\star};\kappa^{\star})-\gamma(\kappa_{t};\kappa^{\star})\big{)}\,\mathrm{d}t\Big{]}\\ &\,\,\,\,-\mathbb{E}_{q}\Big{[}\int_{0}^{T}\lambda^{+}e^{-\kappa^{\star}\psi^{\kappa_{t},+}(Q_{t}^{\psi^{\kappa_{t}};\kappa^{\star}})}\Big{(}\hat{v}^{\psi^{\kappa_{t}}}(Q_{t}^{\psi^{\kappa_{t}};\kappa^{\star}}-1;\kappa^{\star})-\hat{v}^{\psi^{\kappa_{t}}}(Q_{t}^{\psi^{\kappa_{t}};\kappa^{\star}};\kappa^{\star})\Big{)}\bm{1}_{Q_{t}^{\psi^{\kappa_{t}};\kappa^{\star}}>\underline{q}}\\ &\quad\quad+\lambda^{-}e^{-\kappa^{\star}\psi^{\kappa_{t},-}(Q_{t}^{\psi^{\kappa_{t}};\kappa^{\star}})}\Big{(}\hat{v}^{\psi^{\kappa_{t}}}(Q_{t}^{\psi^{\kappa_{t}};\kappa^{\star}}+1;\kappa^{\star})-\hat{v}^{\psi^{\kappa_{t}}}(Q_{t}^{\psi^{\kappa_{t}};\kappa^{\star}};\kappa^{\star})\Big{)}\bm{1}_{Q_{t}^{\psi^{\kappa_{t}};\kappa^{\star}}<\bar{q}}\,\mathrm{d}t\Big{]}\,.\\ \end{split}$
@@
(58) $\begin{split}h(\kappa_{t},q)=\lambda^{+}e^{-\kappa^{\star}\psi^{\kappa_{t},+}(q)}&\big{(}\hat{v}^{\psi^{\kappa_{t}}}(q-1;\kappa^{\star})-\hat{v}^{\psi^{\kappa_{t}}}(q;\kappa^{\star})\big{)}\bm{1}_{q>\underline{q}}\\ &+\lambda^{-}e^{-\kappa^{\star}\psi^{\kappa_{t},-}(q)}\big{(}\hat{v}^{\psi^{\kappa_{t}}}(q+1;\kappa^{\star})-\hat{v}^{\psi^{\kappa_{t}}}(q;\kappa^{\star})\big{)}\bm{1}_{q<\bar{q}}.\end{split}$
@@
where the last inequality comes from Corollary 25. Moreover, $\pi^{\psi^{\kappa_{t}};\kappa^{\star}}_{t}$ is the probability measure evolves under control $\psi^{\kappa_{t}}$ and parameter $\kappa^{\star}$ with $Q_{0}\sim q$. By Lemma 37 (2), there exists a constant $\bar{h}>0$ such that $\left|h(\kappa_{t},q)\right|\leq\bar{h}$ for $\kappa_{t}\in[\underline{K},\bar{K}]$. Therefore,
$\begin{split}\mathcal{R}^{\Psi}(T)&\leq C\mathbb{E}\Big{[}\int_{0}^{T}\left|\kappa_{t}-\kappa^{\star}\right|^{2}\,\mathrm{d}t\Big{]}+\left|\int_{0}^{T}\int_{\Omega}h(\kappa_{t},q)\,\mathrm{d}\pi^{\psi^{\kappa_{t}};\kappa^{\star}}_{t}\,\mathrm{d}t\right|\\ &\leq C\mathbb{E}\Big{[}\int_{0}^{T}\left|\kappa_{t}-\kappa^{\star}\right|^{2}\,\mathrm{d}t\Big{]}+\left|\bar{h}\int_{0}^{T}\int_{\Omega}\frac{h(\kappa_{t},q)}{\bar{h}}(\,\mathrm{d}\pi^{\psi^{\kappa_{t}};\kappa^{\star}}_{t}-\mathrm{d}\pi^{\psi^{\kappa_{t}};\kappa^{\star}})\,\mathrm{d}t\right|\\ &\quad+\left|\int_{0}^{T}\int_{\Omega}h(\kappa_{t},q)\,\mathrm{d}\pi^{\psi^{\kappa_{t}};\kappa^{\star}}\,\mathrm{d}t\right|\\ &\leq C\mathbb{E}\Big{[}\int_{0}^{T}\left|\kappa_{t}-\kappa^{\star}\right|^{2}\,\mathrm{d}t\Big{]}+\left|\bar{h}\int_{0}^{T}\int_{\Omega}\frac{h(\kappa_{t},q)}{\bar{h}}(\,\mathrm{d}\pi^{\psi^{\kappa_{t}};\kappa^{\star}}_{t}-\mathrm{d}\pi^{\psi^{\kappa_{t}};\kappa^{\star}})\,\mathrm{d}t\right|,\end{split}$
where $\pi^{\psi^{\kappa_{t}};\kappa^{\star}}$ is the probability measure of equilibrium distribution under control $\psi^{\kappa_{t}}$ and parameter $\kappa^{\star}$, and the last step uses Proposition 26. Then,
$\begin{split}\mathcal{R}^{\Psi}(T)&\leq C\mathbb{E}\Big{[}\int_{0}^{T}\left|\kappa_{t}-\kappa^{\star}\right|^{2}\,\mathrm{d}t\Big{]}+\bar{h}\int_{0}^{T}\int_{\Omega}\left|\mathrm{d}\pi^{\psi^{\kappa_{t}};\kappa^{\star}}_{t}-\mathrm{d}\pi^{\psi^{\kappa_{t}};\kappa^{\star}}\right|\,\mathrm{d}t\\ &=C\mathbb{E}\Big{[}\int_{0}^{T}\left|\kappa_{t}-\kappa^{\star}\right|^{2}\,\mathrm{d}t\Big{]}+2\bar{h}\int_{0}^{T}\left\|\pi^{\psi^{\kappa_{t}};\kappa^{\star}}_{t}-\pi^{\psi^{\kappa_{t}};\kappa^{\star}}\right\|_{TV}\,\mathrm{d}t,\\ \end{split}$
@@
$\begin{split}\mathcal{R}^{\Psi}(T)&\leq C\mathbb{E}\Big{[}\int_{0}^{T}\left|\kappa_{t}-\kappa^{\star}\right|^{2}\,\mathrm{d}t\Big{]}+2\bar{h}\int_{0}^{T}C(\kappa_{t})\alpha(\kappa_{t})^{t}\,\mathrm{d}t\\ &\leq C\mathbb{E}\Big{[}\int_{0}^{T}\left|\kappa_{t}-\kappa^{\star}\right|^{2}\,\mathrm{d}t\Big{]}+2\bar{h}\bar{C}\int_{0}^{T}\bar{\alpha}^{t}\,\mathrm{d}t\\ &=C\mathbb{E}\Big{[}\int_{0}^{T}\left|\kappa_{t}-\kappa^{\star}\right|^{2}\,\mathrm{d}t\Big{]}+2\bar{h}\bar{C}\frac{e^{T\ln\bar{\alpha}}-1}{\ln\bar{\alpha}},\end{split}$
where $\bar{C}=\sup_{\kappa_{t}}C(\kappa_{t})>0$ and $10<\bar{\alpha}=\sup_{\kappa_{t}}\alpha(\kappa_{t})<$ 1 for $\kappa_{t}\in[\underline{K},\bar{K}]$. The existence of $\bar{C}$ and $\bar{\alpha}$ can be concluded by the following discussion. Since the convergence rate $\alpha$ is actually the second largest eigenvalue of the transition probability matrix $\bm{P}$ [20, 21], which is a continuous function of $\kappa$ on $[\underline{K},\bar{K}]$ by the fact that all eigenvalues for the tridiagonal transition matrix $\bm{P}$ are simple, i.e. the multiplicity is 1. Therefore, there exists $\bar{\alpha}=\sup_{\kappa\in[\underline{K},\bar{K}]}\alpha(\kappa)$. Moreover, $C$ can be given by the norm of the eigenvectors for $\bm{P}$. By [3], if all the eigenvalues of the $\bm{P}$ are simple, the corresponding eigenvectors can be chosen absolutely continuous on $\kappa\in[\underline{K},\bar{K}]$, hence the existence of $\bar{C}$.
Let $C_{1}=C>0,C_{2}=2\bar{h}\bar{C}>0$ and $10<\alpha=\bar{\alpha}<$ 1, we conclude the result. ∎
@@ A.18.1. Proof of Proposition 28
Let $Z_{n}:=f(\delta_{n})(Y_{n}-e^{-\kappa^{\star}\delta_{n}})$. Since $-\|f\|_{\infty}\leq Z_{n}\leq\|f\|_{\infty}$ and since $\mathbb{E}\big{[}Z_{n}\big{|}(\delta_{n})_{n=1}^{N}\big{]}=0$, we get that
@@
$\mathbb{P}^{\star}\left(\sum_{n=1}^{N}Z_{n}>h\bigg{|}(\delta_{n})_{n=1}^{N}\right)\leq\inf_{\lambda\in\mathbb{R}}\exp\left(\frac{N}{2}\lambda^{2}\|f\|_{\infty}^{2}-\lambda h\right)=\exp\left(-\frac{h^{2}}{2N\|f\|_{\infty}^{2}}\right).$
@@
$\mathbb{P}^{\star}\left(\sum_{n=1}^{N}Z_{n}>\|f\|_{\infty}\sqrt{2N\ln\big{(}\tfrac{2}{\epsilon}\big{)}}\bigg{|}(\delta_{n})_{n=1}^{N}\right)\leq\frac{\epsilon}{2}.$
@@ A.18.2. Proof of Proposition 30
$\frac{\mathrm{d}^{2}}{\mathrm{d}\kappa^{2}}\tilde{\ell}_{N}(\kappa)=-$ $\left(\sum_{n=1}^{N}(1-Y_{n})\delta_{n}^{2}\frac{e^{-\kappa\delta_{n}}}{(1-e^{-\kappa\delta_{n}})^{2}}\bm{1}_{\delta_{n}<+\infty}+\delta_{0}^{2}\frac{e^{-\kappa\delta_{0}}}{(1-e^{-\kappa\delta_{0}})^{2}}\right)\bm{1}_{\kappa\leq\bar{K}}$
$-\left(\sum_{n=1}^{N}(1-Y_{n})\delta_{n}^{2}\frac{e^{-\bar{K}\delta_{n}}}{(1-e^{-\bar{K}\delta_{n}})^{2}}\bm{1}_{\delta_{n}<+\infty}+\delta_{0}^{2}\frac{e^{-\bar{K}\delta_{0}}}{(1-e^{-\bar{K}\delta_{0}})^{2}}\right)\bm{1}_{\kappa>\bar{K}}.$
@@
$-\frac{\mathrm{d}^{2}}{\mathrm{d}\kappa^{2}}\tilde{\ell}_{N}(\kappa)$ $\geq\sum_{n=1}^{N}(1-Y_{n})\delta_{n}^{2}\left(\frac{e^{-\kappa\delta_{n}}}{(1-e^{-\kappa\delta_{n}})^{2}}\bm{1}_{\kappa\leq\bar{K}}+\frac{e^{-\bar{K}\delta_{n}}}{(1-e^{-\bar{K}\delta_{n}})^{2}}\bm{1}_{\kappa>\bar{K}}\right)\bm{1}_{\delta_{n}<+\infty}$
$\geq\sum_{n=1}^{N}(1-Y_{n})\delta_{n}^{2}\left(\frac{e^{-\bar{K}\delta_{n}}}{(1-e^{-\bar{K}\delta_{n}})^{2}}\right)\bm{1}_{\delta_{n}<+\infty}.$
Hence, for any policy $(\delta_{n})_{n=1}^{\infty}$ taking values in $[\underline{\delta},\bar{\delta}]\cup\{+\infty\}$, it holds that for any $\kappa\geq 0$,
@@
By Proposition 28, it holds that with $\mathbb{P}^{\star}$-probability at least $1-\epsilon$
$\left|\sum_{n=1}^{N}(Y_{n}-e^{-\kappa^{\star}\delta_{n}})\left(\frac{\underline{\delta}^{2}e^{-\bar{K}\bar{\delta}}}{(1-e^{-\bar{K}\bar{\delta}})^{2}}\right)\right|\leq\left(\frac{\underline{\delta}^{2}e^{-\bar{K}\bar{\delta}}}{(1-e^{-\bar{K}\bar{\delta}})^{2}}\right)\sqrt{2N\ln\left(\tfrac{2}{\epsilon}\right)}.$
@@
$\inf_{\kappa\geq 0}\Big{(}-\frac{\mathrm{d}^{2}}{\mathrm{d}\kappa^{2}}\tilde{\ell}_{N}(\kappa)\Big{)}\geq\sum_{n=1}^{N}(1-Y_{n})\left(\frac{\underline{\delta}^{2}e^{-\bar{K}\bar{\delta}}}{(1-e^{-\bar{K}\bar{\delta}})^{2}}\right)$
$\qquad\geq\sum_{n=1}^{N}(1-e^{-\kappa^{\star}\delta_{n}})\left(\frac{\underline{\delta}^{2}e^{-\bar{K}\bar{\delta}}}{(1-e^{-\bar{K}\bar{\delta}})^{2}}\right)-\left|\sum_{n=1}^{N}(Y_{n}-e^{-\kappa^{\star}\delta_{n}})\left(\frac{\underline{\delta}^{2}e^{-\bar{K}\bar{\delta}}}{(1-e^{-\bar{K}\bar{\delta}})^{2}}\right)\right|$
$\qquad\geq\left(\frac{\underline{\delta}^{2}e^{-\bar{K}\bar{\delta}}(1-e^{-\kappa^{\star}\underline{\delta}})}{(1-e^{-\bar{K}\bar{\delta}})^{2}}\right)N-\left(\frac{\underline{\delta}^{2}e^{-\bar{K}\bar{\delta}}}{(1-e^{-\bar{K}\bar{\delta}})^{2}}\right)\sqrt{2N\ln\left(\tfrac{2}{\epsilon}\right)}.$

### v2

@@ A.17. Proof of Theorem 30
$\begin{split}\lambda^{+}\psi^{\kappa,+}(q)e^{-\kappa^{\ast}\psi^{\kappa,+}(q)}+\lambda^{-}\psi^{\kappa,-}(q)&e^{-\kappa^{\ast}\psi^{\kappa,-}(q)}-\phi q^{2}=\\ \gamma(\kappa;\kappa^{\ast})&+\lambda^{+}e^{-\kappa^{\ast}\psi^{\kappa,+}}\big{(}\hat{v}^{\psi^{\kappa}}(q-1;\kappa^{\ast})-\hat{v}^{\psi^{\kappa}}(q;\kappa^{\ast})\big{)}\mathds{1}_{q>\underline{q}}\\ &+\lambda^{-}e^{-\kappa^{\ast}\psi^{\kappa,-}}\big{(}\hat{v}^{\psi^{\kappa}}(q+1;\kappa^{\ast})-\hat{v}^{\psi^{\kappa}}(q;\kappa^{\ast})\big{)}\mathds{1}_{q<\bar{q}},\end{split}$
where we ignore the indicator functions in the first line since $\psi^{\kappa,\pm}(q)e^{-\kappa^{\ast}\psi^{\kappa,\pm}(q)}=0$ for $q=\bar{q},\underline{q}$, respectively. Also, by definition, we know that
$f(t,q,\delta^{\pm};\kappa^{\ast})=\lambda^{+}\delta^{+}e^{-\kappa^{\ast}\delta^{+}}+\lambda^{-}\delta^{-}e^{-\kappa^{\ast}\delta^{-}}-\phi q^{2}.$
@@
(61) $\begin{split}&\mathcal{R}^{\Psi}(T)=\gamma(\kappa^{\ast})T-\mathbb{E}_{q}\Big{[}\int_{0}^{T}f(t,Q_{t}^{\psi^{\kappa_{t}};\kappa^{\ast}},\psi^{\kappa_{t}};\kappa^{\ast})\,\mathrm{d}t\Big{]}\\ &=\mathbb{E}\Big{[}\int_{0}^{T}\big{(}\gamma(\kappa^{\ast};\kappa^{\ast})-\gamma(\kappa_{t};\kappa^{\ast})\big{)}\,\mathrm{d}t\Big{]}\\ &\,\,\,\,-\mathbb{E}_{q}\Big{[}\int_{0}^{T}\lambda^{+}e^{-\kappa^{\ast}\psi^{\kappa_{t},+}(Q_{t}^{\psi^{\kappa_{t}};\kappa^{\ast}})}\Big{(}\hat{v}^{\psi^{\kappa_{t}}}(Q_{t}^{\psi^{\kappa_{t}};\kappa^{\ast}}-1;\kappa^{\ast})-\hat{v}^{\psi^{\kappa_{t}}}(Q_{t}^{\psi^{\kappa_{t}};\kappa^{\ast}};\kappa^{\ast})\Big{)}\mathds{1}_{Q_{t}^{\psi^{\kappa_{t}};\kappa^{\ast}}>\underline{q}}\\ &\quad\quad+\lambda^{-}e^{-\kappa^{\ast}\psi^{\kappa_{t},-}(Q_{t}^{\psi^{\kappa_{t}};\kappa^{\ast}})}\Big{(}\hat{v}^{\psi^{\kappa_{t}}}(Q_{t}^{\psi^{\kappa_{t}};\kappa^{\ast}}+1;\kappa^{\ast})-\hat{v}^{\psi^{\kappa_{t}}}(Q_{t}^{\psi^{\kappa_{t}};\kappa^{\ast}};\kappa^{\ast})\Big{)}\mathds{1}_{Q_{t}^{\psi^{\kappa_{t}};\kappa^{\ast}}<\bar{q}}\,\mathrm{d}t\Big{]}\,.\\ \end{split}$
@@
(62) $\begin{split}h(\kappa_{t},q)=\lambda^{+}e^{-\kappa^{\ast}\psi^{\kappa_{t},+}(q)}&\big{(}\hat{v}^{\psi^{\kappa_{t}}}(q-1;\kappa^{\ast})-\hat{v}^{\psi^{\kappa_{t}}}(q;\kappa^{\ast})\big{)}\mathds{1}_{q>\underline{q}}\\ &+\lambda^{-}e^{-\kappa^{\ast}\psi^{\kappa_{t},-}(q)}\big{(}\hat{v}^{\psi^{\kappa_{t}}}(q+1;\kappa^{\ast})-\hat{v}^{\psi^{\kappa_{t}}}(q;\kappa^{\ast})\big{)}\mathds{1}_{q<\bar{q}}.\end{split}$
@@
where the last inequality comes from Corollary 24. Moreover, $\pi^{\psi^{\kappa_{t}};\kappa^{\ast}}_{t}$ is the probability measure evolves under control $\psi^{\kappa_{t}}$ and parameter $\kappa^{\ast}$ with $Q_{0}\sim q$. By Lemma 39 (2), there exists a constant $\bar{h}>0$ such that $\left|h(\kappa_{t},q)\right|\leq\bar{h}$ for $\kappa_{t}\in[\underline{K},\bar{K}]$. Therefore,
$\begin{split}\mathcal{R}^{\Psi}(T)&\leq C\mathbb{E}\Big{[}\int_{0}^{T}\left|\kappa_{t}-\kappa^{\ast}\right|^{2}\,\mathrm{d}t\Big{]}+\left|\int_{0}^{T}\int_{\Omega}h(\kappa_{t},q)\,\mathrm{d}\pi^{\psi^{\kappa_{t}};\kappa^{\ast}}_{t}\,\mathrm{d}t\right|\\ &\leq C\mathbb{E}\Big{[}\int_{0}^{T}\left|\kappa_{t}-\kappa^{\ast}\right|^{2}\,\mathrm{d}t\Big{]}+\left|\bar{h}\int_{0}^{T}\int_{\Omega}\frac{h(\kappa_{t},q)}{\bar{h}}(\,\mathrm{d}\pi^{\psi^{\kappa_{t}};\kappa^{\ast}}_{t}-\mathrm{d}\pi^{\psi^{\kappa_{t}};\kappa^{\ast}})\,\mathrm{d}t\right|\\ &\quad+\left|\int_{0}^{T}\int_{\Omega}h(\kappa_{t},q)\,\mathrm{d}\pi^{\psi^{\kappa_{t}};\kappa^{\ast}}\,\mathrm{d}t\right|\\ &\leq C\mathbb{E}\Big{[}\int_{0}^{T}\left|\kappa_{t}-\kappa^{\ast}\right|^{2}\,\mathrm{d}t\Big{]}+\left|\bar{h}\int_{0}^{T}\int_{\Omega}\frac{h(\kappa_{t},q)}{\bar{h}}(\,\mathrm{d}\pi^{\psi^{\kappa_{t}};\kappa^{\ast}}_{t}-\mathrm{d}\pi^{\psi^{\kappa_{t}};\kappa^{\ast}})\,\mathrm{d}t\right|,\end{split}$
where $\pi^{\psi^{\kappa_{t}};\kappa^{\ast}}$ is the probability measure of equilibrium distribution under control $\psi^{\kappa_{t}}$ and parameter $\kappa^{\ast}$, and the last step uses Proposition 29. Then,
$\begin{split}\mathcal{R}^{\Psi}(T)&\leq C\mathbb{E}\Big{[}\int_{0}^{T}\left|\kappa_{t}-\kappa^{\ast}\right|^{2}\,\mathrm{d}t\Big{]}+\bar{h}\int_{0}^{T}\int_{\Omega}\left|\mathrm{d}\pi^{\psi^{\kappa_{t}};\kappa^{\ast}}_{t}-\mathrm{d}\pi^{\psi^{\kappa_{t}};\kappa^{\ast}}\right|\,\mathrm{d}t\\ &=C\mathbb{E}\Big{[}\int_{0}^{T}\left|\kappa_{t}-\kappa^{\ast}\right|^{2}\,\mathrm{d}t\Big{]}+2\bar{h}\int_{0}^{T}\left\|\pi^{\psi^{\kappa_{t}};\kappa^{\ast}}_{t}-\pi^{\psi^{\kappa_{t}};\kappa^{\ast}}\right\|_{TV}\,\mathrm{d}t,\\ \end{split}$
@@
$\begin{split}\mathcal{R}^{\Psi}(T)&\leq C\mathbb{E}\Big{[}\int_{0}^{T}\left|\kappa_{t}-\kappa^{\ast}\right|^{2}\,\mathrm{d}t\Big{]}+2\bar{h}\int_{0}^{T}C(\kappa_{t})\alpha(\kappa_{t})^{t}\,\mathrm{d}t\\ &\leq C\mathbb{E}\Big{[}\int_{0}^{T}\left|\kappa_{t}-\kappa^{\ast}\right|^{2}\,\mathrm{d}t\Big{]}+2\bar{h}\bar{C}\int_{0}^{T}\bar{\alpha}^{t}\,\mathrm{d}t\\ &=C\mathbb{E}\Big{[}\int_{0}^{T}\left|\kappa_{t}-\kappa^{\ast}\right|^{2}\,\mathrm{d}t\Big{]}+2\bar{h}\bar{C}\frac{1-e^{T\ln\bar{\alpha}}}{\ln(\bar{\alpha}^{-1})}\\ &\leq C\mathbb{E}\Big{[}\int_{0}^{T}\left|\kappa_{t}-\kappa^{\ast}\right|^{2}\,\mathrm{d}t\Big{]}+\frac{2\bar{h}\bar{C}}{\ln(\alpha^{-1})}\,,\end{split}$
where $\bar{C}=\sup_{\kappa_{t}}C(\kappa_{t})>0$ and $10<\bar{\alpha}=\sup_{\kappa_{t}}\alpha(\kappa_{t})<$ 1 for $\kappa_{t}\in[\underline{K},\bar{K}]$. The existence of $\bar{C}$ and $\bar{\alpha}$ can be concluded by the following discussion. As shown in Appendix A.15, the state dynamics can be represented by a continuous-time Markov chain. Therefore, the convergence rate $\alpha$ is bounded by the exponential of the second largest eigenvalue of the transition rate matrix $\bm{Q}$ by the Kolmogorov forward equation [6, Chapter 6.6]. Since the transition rate matrix $\bm{Q}$ (60) is tridiagonal, its eigenvalues are simple, i.e. the multiplicity is 1, implying that the eigenvalues are continuous with respect to $\kappa$. Therefore, there exists $\bar{\alpha}=\sup_{\kappa\in[\underline{K},\bar{K}]}\alpha(\kappa)$. Moreover, $C$ can be given by the norm of the eigenvectors for $\bm{Q}$. By [3], if all the eigenvalues of the $\bm{Q}$ are simple, the corresponding eigenvectors can be chosen absolutely continuous on $\kappa\in[\underline{K},\bar{K}]$, hence the existence of $\bar{C}$. Let $C_{1}=C>0,C_{2}=2\bar{h}\bar{C}>0$ and $10<\alpha=\bar{\alpha}<$ 1, we conclude the result. ∎
@@ A.18.1. Proof of Proposition 31
Let $Z_{n}:=f(\delta_{n})(Y_{n}-e^{-\kappa^{\ast}\delta_{n}})$. Since $-\|f\|_{\infty}\leq Z_{n}\leq\|f\|_{\infty}$ and since $\mathbb{E}\big{[}Z_{n}\big{|}(\delta_{n})_{n=1}^{N}\big{]}=0$, we get that
@@
$\mathbb{P}^{\ast}\left(\sum_{n=1}^{N}Z_{n}>h\bigg{|}(\delta_{n})_{n=1}^{N}\right)\leq\inf_{\lambda\in\mathbb{R}}\exp\left(\frac{N}{2}\lambda^{2}\|f\|_{\infty}^{2}-\lambda h\right)=\exp\left(-\frac{h^{2}}{2N\|f\|_{\infty}^{2}}\right).$
@@
$\mathbb{P}^{\ast}\left(\sum_{n=1}^{N}Z_{n}>\|f\|_{\infty}\sqrt{2N\ln\big{(}\tfrac{2}{\varepsilon}\big{)}}\bigg{|}(\delta_{n})_{n=1}^{N}\right)\leq\frac{\varepsilon}{2}.$
@@ A.18.2. Proof of Proposition 32
$\frac{\mathrm{d}^{2}}{\mathrm{d}\kappa^{2}}\tilde{\ell}_{N}(\kappa)=-$ $\left(\sum_{n=1}^{N}(1-Y_{n})\delta_{n}^{2}\frac{e^{-\kappa\delta_{n}}}{(1-e^{-\kappa\delta_{n}})^{2}}+\delta_{0}^{2}\frac{e^{-\kappa\delta_{0}}}{(1-e^{-\kappa\delta_{0}})^{2}}\right)\mathds{1}_{\kappa\leq\bar{K}}$
$\qquad-\left(\sum_{n=1}^{N}(1-Y_{n})\delta_{n}^{2}\frac{e^{-\bar{K}\delta_{n}}}{(1-e^{-\bar{K}\delta_{n}})^{2}}+\delta_{0}^{2}\frac{e^{-\bar{K}\delta_{0}}}{(1-e^{-\bar{K}\delta_{0}})^{2}}\right)\mathds{1}_{\kappa>\bar{K}}.$
@@
$-\frac{\mathrm{d}^{2}}{\mathrm{d}\kappa^{2}}\tilde{\ell}_{N}(\kappa)$ $\geq\sum_{n=1}^{N}(1-Y_{n})\delta_{n}^{2}\left(\frac{e^{-\kappa\delta_{n}}}{(1-e^{-\kappa\delta_{n}})^{2}}\mathds{1}_{\kappa\leq\bar{K}}+\frac{e^{-\bar{K}\delta_{n}}}{(1-e^{-\bar{K}\delta_{n}})^{2}}\mathds{1}_{\kappa>\bar{K}}\right)$
$\geq\sum_{n=1}^{N}(1-Y_{n})\delta_{n}^{2}\left(\frac{e^{-\bar{K}\delta_{n}}}{(1-e^{-\bar{K}\delta_{n}})^{2}}\right)\,.$
Hence, for any policy $(\delta_{n})_{n=1}^{\infty}$ taking values in $[\underline{\delta},\bar{\delta}]\cup\{+\infty\}$, it holds that for any $\kappa>0$,
@@
By Proposition 31, it holds that with $\mathbb{P}^{\ast}$-probability at least $1-\varepsilon$
$\left|\sum_{n=1}^{N}(Y_{n}-e^{-\kappa^{\ast}\delta_{n}})\left(\frac{\underline{\delta}^{2}e^{-\bar{K}\bar{\delta}}}{(1-e^{-\bar{K}\bar{\delta}})^{2}}\right)\right|\leq\left(\frac{\underline{\delta}^{2}e^{-\bar{K}\bar{\delta}}}{(1-e^{-\bar{K}\bar{\delta}})^{2}}\right)\sqrt{2N\ln\left(\tfrac{2}{\varepsilon}\right)}.$
@@
$\inf_{\kappa>0}\Big{(}-\frac{\mathrm{d}^{2}}{\mathrm{d}\kappa^{2}}\tilde{\ell}_{N}(\kappa)\Big{)}\geq\sum_{n=1}^{N}(1-Y_{n})\left(\frac{\underline{\delta}^{2}e^{-\bar{K}\bar{\delta}}}{(1-e^{-\bar{K}\bar{\delta}})^{2}}\right)$
$\qquad\geq\sum_{n=1}^{N}(1-e^{-\kappa^{\ast}\delta_{n}})\left(\frac{\underline{\delta}^{2}e^{-\bar{K}\bar{\delta}}}{(1-e^{-\bar{K}\bar{\delta}})^{2}}\right)-\left|\sum_{n=1}^{N}(Y_{n}-e^{-\kappa^{\ast}\delta_{n}})\left(\frac{\underline{\delta}^{2}e^{-\bar{K}\bar{\delta}}}{(1-e^{-\bar{K}\bar{\delta}})^{2}}\right)\right|$
$\qquad\geq\left(\frac{\underline{\delta}^{2}e^{-\bar{K}\bar{\delta}}(1-e^{-\underline{K}\,\underline{\delta}})}{(1-e^{-\bar{K}\bar{\delta}})^{2}}\right)N-\left(\frac{\underline{\delta}^{2}e^{-\bar{K}\bar{\delta}}}{(1-e^{-\bar{K}\bar{\delta}})^{2}}\right)\sqrt{2N\ln\left(\tfrac{2}{\varepsilon}\right)}.$